# Ultralight Polarity-Split Neuromorphic SNN for Event-Stream Super-Resolution

### v1

@@ 1 Introduction
Event cameras, also known as neuromorphic cameras or dynamic vision sensors (DVS), are asynchronous sensors that respond to changes in scene brightness. When the brightness change at a given pixel exceeds a certain threshold, an event is triggered and recorded as a tuple $(x_{k},y_{k},t_{k},p_{k})$, where $(x_{k},y_{k})$ denotes the spatial coordinates of the pixel, $t_{k}$ is the precise timestamp, and $p_{k}\in\{+1,-1\}$ represents the polarity of brightness change [1, 2]. The continuous sequence of $N$ events forms an event stream, which can be represented as:
@@
Due to this asynchronous and sparse sensing mechanism, event streams exhibit several unique advantages over traditional frame-based RGB data, including high temporal resolution, high dynamic range (HDR), low latency, and negligible motion blur [1, 3]. These properties make them especially promising for applications in 3D reconstruction [4], high-speed robotics [5], AR/VR [6], and autonomous driving [7].
@@
The spatial resolution of most commercially available event cameras ($\leq 640\times 480$) remains significantly lower than that of frame-based cameras [8]. Although high-resolution event sensors such as the Sony IMX646 ($1280\times 720$) have been developed [8], achieving higher resolution at the hardware level introduces increased power consumption and cost [9]. Moreover, recent studies suggest that the motivation for developing higher-resolution event cameras may be limited [10]. In extreme conditions such as low-light scenes or high-speed motion, high-resolution event cameras may perform worse, since they elevate per-pixel event rates, which in turn amplify temporal noise [10]. On the other hand, more studies show that under standard conditions, low-resolution event data limits fine-grained perception and downstream performance, while inputting high-resolution events can improve tasks like object recognition and image/video reconstruction [32, 9, 12, 31].
@@
Existing event super-resolution methods can be broadly categorized into two types, referring to Figure 2 and Section 2. Event-to-frame super-resolution converts event streams into frame-based representations such as event stacks or event count maps [29, 30, 31], and then applies reconstruction algorithms or networks similar to those used in image super-resolution. The frame-based output is then uniformly or randomly sampled on the temporal dimension to produce high-resolution event streams. As many downstream vision tasks still require frame-based inputs to interface with processing modules like CNNs, the loss of temporal precision caused during this process is often considered acceptable. The second type is event-to-event stream-based super-resolution, which directly reconstructs high-resolution event streams without any temporal sampling [32, 16]. It preserves the asynchronous and temporally precise nature of event data, while recovering such temporally asynchronous events requires specialized networks such as Spiking Neural Networks (SNNs) [32].
@@
More importantly, an ideal event super-resolution module should be lightweight, energy-efficient, and capable of real-time processing. It is impractical to allocate a high-end chip solely for running a heavy model on an event camera, and it is also undesirable to dedicate substantial computational resources to a super-resolution module within a vision processing pipeline as a step of preprocessing. However, recent methods have made this module still heavy [16, 12, 31].
@@ 2.1 Event-to-Frame Super-Resolution
Before 2020, event-to-frame super-resolution was mainly based on mathematical modeling. Li et al. were the first to address this task [29]. They modeled the event stream at each pixel as a non-homogeneous Poisson process, using event count maps to recover the number of events in the spatial domain and spatiotemporal filters to estimate the event rate function in the temporal domain. Later, Wang et al. proposed the Guided Event Filtering method [17], which jointly filtered frame images and event data to improve spatial denoising in event super-resolution.
Subsequently, deep learning-based methods were developed. In 2021, Duan et al. introduced EventZoom [30], a method based on a 3D U-Net architecture with an event-to-image module to leverage high-resolution image features, achieving better results than previous non-learning approaches. In 2022, Weng et al. proposed RecEvSR [9], which used an RNN with a spatial attention mechanism to enhance detail in key areas, showing strong performance for high upsampling rates.
More recent works have explored separating polarities and feeding them into different branches of a network, with interaction modules to exchange information. In 2024, Huang et al. proposed BMCNet [12], a bilateral mining and complementary fusion framework that extracts positive and negative polarity features separately, then fuses them to improve detail recovery. Liang et al. proposed RMFNet [31], which combines a dual-branch architecture with a recursive structure to enhance temporal modeling across frames and further improve performance.
@@ 2.2 Event-to-Event Super-Resolution
In 2021, Li et al. introduced an innovative approach using SNNs for the event super-resolution task [32]. SNNs can directly process raw event streams and output high-resolution asynchronous event streams, maintaining the original temporal characteristics. In addition, they designed spatiotemporal constraints to enable the network to learn both spatial and temporal distributions of events. This method serves as the primary baseline for our work.
Figure 3: Architecture of key modules in our proposed event super-resolution method. Left: SRM-based spiking neuron structure and spike triggering mechanism. Center: Main architecture of our method. Right: Composition of the Learnable Spatio-temporal Polarity-aware Loss function. (a) Dual-layer EventSR Network. (b) Dual-Forward Polarity-Split Event Encoding Strategy. (c) Ultra-lightweight Polarity-Split EventSR Network.
@@ 3.1 SRM-based Spiking Neuron Encoding
We adopt a Spiking Neural Network (SNN) based on the Spike Response Model (SRM) [39] to address the task of event stream super-resolution. Unlike conventional artificial neurons that rely on continuous and differentiable activation functions, SRM-based spiking neurons encode and transmit information through temporal spike trains, simulating biological neural behavior.
@@ 3.2 Dual-layer EventSR SNN
Most previous works tend to increase the complexity of neural architectures for event stream super-resolution to achieve better performance [12, 31, 16]. However, the baseline work has shown that for SNN-based event stream super-resolution, deeper networks may actually degrade performance [32], which supports our motivation to further simplify the network design.
We propose a compact yet effective SNN-based event stream super-resolution network, simply named Dual-layer SNN in this paper. As shown in Figure 3a and Table 1, this network consists of two primary spiking convolution layers. The input and output of the network contain two channels, representing positive and negative event streams. The first layer extracts local spatiotemporal spike patterns and generates intermediate spike responses. The second layer upsamples the output of the first layer back to the original resolution. As shown in Figure 4, it also incorporates an upsampling bilinear-interpolated PSP bypass to preserve fine details from the low-resolution input for final reconstruction.
@@ 3.3 Dual-Forward Polarity-split Event Encoding
Recent studies have proposed using dual-branch networks to separately process positive and negative events [12, 31], demonstrating that handling polarities separately leads to improved performance on event super-resolution. However, the additional network branches significantly increase the model’s weight.
@@ 3.5 Learnable Spatio-temporal Polarity-aware Loss
Inspired by previous work that performs the spatio-temporal learning [32] and weighted uncertainty loss [19], we extend the spatio-temporal only loss with polarity awareness and learnable adaptive weighting, namely Learnable Spatio-temporal Polarity-aware Loss (LearnSTPLoss), simply marked as "w/L" in this paper. Given the predicted spike output $e_{\text{out}}$ and the ground truth spike stream $e_{\text{gt}}$, we have:
@@
Preserving time, space, and polarity information is essential to the event stream super-resolution task. When using a combined loss function, previous methods often rely on manually set hyperparameters to weight each component [32, 16]. However, such manual tuning lacks adaptability across different datasets or model architectures. To address this issue, we introduce a learnable weighting mechanism based on the logarithm of variance, enabling the weights of each loss term to be jointly balanced and optimized with the network parameters during training. This learnable loss formulation is expected to improve super-resolution accuracy without compromising the model’s lightweight design.
@@ 4.1 Experimental Settings
We evaluate our method on five datasets, as summarized in Table 3. The HR event streams correspond to the original event data in these datasets. To generate the LR event streams, we perform spatial downsampling by merging events within each 2×2 region using a stride of 2.
Dataset Train/Test Sensor Batch Size
N-MNIST [20] 60k / 10k ATIS 64
CIFAR10-DVS [21] 8.5k / 1.5k DVS128 8
ASL-DVS [22] 80.6k / 20.2k DVS128 32
EventCameraD. [23] 10.1k / 1.6k DAVIS240C 64
EventNFS-real [30] 74.2k / 10.5k DAVIS346 64
Table 3: Datasets Information and Batch Size in Training
@@
(a) Dual-Layer SNN .2621 .1292 .2280 97.34 .1781 .1002 .1473 69.92 .2250 .1048 .1975 99.39 .3165 .2593 .1697 89.06 .3180 .2160 .2329 87.46
(b) Dual-Layer SNN † .2607 .1277 .2272 97.87 .1772 .1017 .1451 70.17 .2250 .1038 .1982 99.42 .3130 .2626 .1675 89.82 .3105 .2148 .2237 87.65
(c) Ultralight SNN .2614 .1289 .2274 95.63 .1754 .0992 .1446 68.28 .2247 .1069 .1961 98.87 .3128 .2626 .1674 89.85 .3113 .2123 .2270 89.35
(d) Ultralight SNN † .2611 .1283 .2274 95.74 .1747 .0994 .1437 68.78 .2236 .1067 .1948 98.93 .3117 .2592 .1699 89.03 .3061 .2118 .2202 89.96
Table 4: Comparison of our methods and baseline. LearnSTPLoss is marked as †. (The best; The second best.)
Implementation Details.
We implement our SNN architecture using SlayerSNN [24], with a discretized simulation step size of 1 millisecond to update membrane potentials and determine spike firing. We train our model for 30 epochs using the Adam optimizer with an initial learning rate of 0.1. Table 3 provides the batch size for training each of the datasets. All experiments are conducted on an NVIDIA RTX 5090 GPU.
4.2 Evaluation Metrics
Evaluation of Event Stream.
To quantitatively evaluate the performance of event stream super-resolution, referring to our baseline [32], we implement a Root Mean Squared Error (RMSE) metric that considers both temporal and spatial precision. The predicted and ground-truth event streams are represented as four-column arrays $(t,x,y,p)$ and discretized into 4D voxel tensors of shape (C=2,H,W,T)(C{=}2,H,W,T), where the two channels correspond to positive and negative polarities.
The temporal consistency error $\mathrm{MSE}_{\mathrm{T}}$ is computed by summing the squared differences of the voxel tensors across all spatiotemporal locations [32]. To measure the spatial consistency error $\mathrm{MSE}_{\mathrm{S}}$, we divide the event stream into non-overlapping time bins (50 ms each), compute the Peri-Stimulus Time Histogram (PSTH) for each pixel, and calculate the squared difference between predicted and ground-truth PSTHs over all bins and pixels [32].
@@ 4.2 Evaluation Metrics
$\mathrm{RMSE}_{\mathrm{ST}}=\sqrt{\frac{1}{(T_{1}-T_{0})\cdot N_{p}}\left(\mathrm{MSE}_{\mathrm{T}}+\mathrm{MSE}_{\mathrm{S}}\right)},$ (15)
where $N_{p}$ is the number of active pixels and $(T_{1}-T_{0})$ is the total time span of the event stream.
@@ 4.3 Comparative Results
Method Params FLOPs (G) Time (ms)
@@
Dual-Layer SNN w/L (ours) 464 (0.464K) 0.53 1.49
Ultralight SNN w/L (ours) 232 (0.232K) 0.58 1.55†
Baseline [32] 1040 (1.04K) 1.10 1.62
@@
SRFBN-esr [25] 2.1M (2100K) 39.5 37.3
RLSP-esr [26] 1.2M (1200K) 23.1
RSTT-esr [27] 3.8M (3800K) 22.3 61.4
@@
EventZoom [30] 11.5M (11500K) 65.3 17.4
RecEvSR [9] 1.8M (1800K) 2.80 13.2
BMCNET [12] 2.6M (2600K) 35.35
RMFNET [31] 3.0M (3000K) 7.0
Table 5: Comparison of model size, computational cost, and inference time on EventNFS dataset. †: Ultralight SNN uses concurrent dual forwards. Parts of the data are from [12].
@@
Referring Table 6, we evaluate our method on the image reconstruction task using the Event Camera Dataset and E2VID [28]. The results show that our method improves SSIM by 7% and reduces MSE by 4.48% compared to the baseline. Figure 8 visualizes the reconstruction, where zoomed-up regions reveal that our method preserves clearer edge details. Additionally, we conduct classification tasks on N-MNIST, ASL-DVS, and CIFAR10-DVS datasets. Our method consistently improves the performance (See Appendix). These results also demonstrate that higher-resolution event data leads to better performance on these tasks, further validating the significance of super-resolving event data.
Metric LR HR-GT [29] Baseline (a) (b)
@@
Table 6: Results of Image Reconstruction on Event Camera Dataset. (a) Dual-Layer SNN w/L (b) Ultralight SNN w/L
@@
Since event-to-frame methods evaluate only the spatial quality of their frame-based HR outputs, our method is not directly comparable at the event stream level. In the appendix, we provide further analytical comparisons in terms of spatiotemporal quality. Additionally, we deploy a DVXplorer-Lite event camera in a mobile scenario to collect real-world data and evaluate our method, demonstrating its robustness in performing 2× super-resolution on raw event streams. (See Appendix)

### v2

@@ 1 Introduction
Event cameras, also known as neuromorphic cameras or dynamic vision sensors (DVS), are asynchronous sensors that respond to changes in scene brightness. When the brightness change at a given pixel exceeds a certain threshold, an event is triggered and recorded as a tuple $(x_{k},y_{k},t_{k},p_{k})$, where $(x_{k},y_{k})$ denotes the spatial coordinates of the pixel, $t_{k}$ is the precise timestamp, and $p_{k}\in\{+1,-1\}$ represents the polarity of brightness change (Intro3; Survey). The continuous sequence of $N$ events forms an event stream, which can be represented as:
@@
Due to this asynchronous and sparse sensing mechanism, event streams exhibit several unique advantages over traditional frame-based RGB data, including high temporal resolution, high dynamic range (HDR), low latency, and negligible motion blur (Intro3; Survey12). These properties make them especially promising for applications in 3D reconstruction (Intro2), high-speed robotics (Survey23), AR/VR (Intro4), and autonomous driving (Survey22).
@@
The spatial resolution of most commercially available event cameras ($\leq 640\times 480$) remains significantly lower than that of frame-based cameras (SurveyEventCameraModel). Although high-resolution event sensors such as the Sony IMX646 ($1280\times 720$) have been developed (SurveyEventCameraModel), achieving higher resolution at the hardware level introduces increased power consumption and cost (EF4). Moreover, recent studies suggest that the motivation for developing higher-resolution event cameras may be limited (Intro1). In extreme conditions such as low-light scenes or high-speed motion, high-resolution event cameras may perform worse, since they elevate per-pixel event rates, which in turn amplify temporal noise (Intro1). On the other hand, more studies show that under standard conditions, low-resolution event data limits fine-grained perception and downstream performance, while inputting high-resolution events can improve tasks like object recognition and image/video reconstruction (ES1; EF4; EF5; EF6).
@@
Existing event super-resolution methods can be broadly categorized into two types, referring to Figure 2 and Section 2. Event-to-frame super-resolution converts event streams into frame-based representations such as event stacks or event count maps (EF1; EF3; EF6), and then applies reconstruction algorithms or networks similar to those used in image super-resolution. The frame-based output is then uniformly or randomly sampled on the temporal dimension to produce high-resolution event streams. As many downstream vision tasks still require frame-based inputs to interface with processing modules like CNNs, the loss of temporal precision caused during this process is often considered acceptable. The second type is event-to-event stream-based super-resolution, which directly reconstructs high-resolution event streams without any temporal sampling (ES1). It preserves the asynchronous and temporally precise nature of event data, while recovering such temporally asynchronous events requires specialized networks such as Spiking Neural Networks (SNNs) (ES1).
@@
More importantly, an ideal event super-resolution module should be lightweight, energy-efficient, and capable of real-time processing. It is impractical to allocate a high-end chip solely for running a heavy model on an event camera, and it is also undesirable to dedicate substantial computational resources to a super-resolution module within a vision processing pipeline as a step of preprocessing. However, recent methods have made this module still heavy (EF5; EF6).
@@ 2.1 Event-to-Frame Super-Resolution
Before 2020, event-to-frame super-resolution was mainly based on mathematical modeling. Li et al. were the first to address this task (EF1). They modeled the event stream at each pixel as a non-homogeneous Poisson process, using event count maps to recover the number of events in the spatial domain and spatiotemporal filters to estimate the event rate function in the temporal domain. Later, Wang et al. proposed the Guided Event Filtering method (EF2), which jointly filtered frame images and event data to improve spatial denoising in event super-resolution.
Subsequently, deep learning-based methods were developed. In 2021, Duan et al. introduced EventZoom (EF3), a method based on a 3D U-Net architecture with an event-to-image module to leverage high-resolution image features, achieving better results than previous non-learning approaches. In 2022, Weng et al. proposed RecEvSR (EF4), which used an RNN with a spatial attention mechanism to enhance detail in key areas, showing strong performance for high upsampling rates.
More recent works have explored separating polarities and feeding them into different branches of a network, with interaction modules to exchange information. In 2024, Huang et al. proposed BMCNet (EF5), a bilateral mining and complementary fusion framework that extracts positive and negative polarity features separately, then fuses them to improve detail recovery. Liang et al. proposed RMFNet (EF6), which combines a dual-branch architecture with a recursive structure to enhance temporal modeling across frames and further improve performance.
Figure 3: Architecture of key modules in our proposed event super-resolution method. Left: Main architecture of our method. Right: Composition of the Learnable Spatio-temporal Polarity-aware Loss function. (a) Dual-layer EventSR Network. (b) Dual-Forward Polarity-Split Event Encoding Strategy. (c) Ultra-lightweight Polarity-Split EventSR Network.
Figure 4: SRM-based spiking neuron structure and spike triggering mechanism.
@@ 2.2 Event-to-Event Super-Resolution
In 2021, Li et al. introduced an innovative approach using SNNs for the event super-resolution task (ES1). SNNs can directly process raw event streams and output high-resolution asynchronous event streams, maintaining the original temporal characteristics. In addition, they designed spatiotemporal constraints to enable the network to learn both spatial and temporal distributions of events. This method serves as the primary baseline for our work.
@@ 3.1 SRM-based Spiking Neuron Encoding
We adopt a Spiking Neural Network (SNN) based on the Spike Response Model (SRM) (SRM) to address the task of event stream super-resolution. Unlike conventional artificial neurons that rely on continuous and differentiable activation functions, SRM-based spiking neurons encode and transmit information through temporal spike trains, simulating biological neural behavior, as shown in Figure 4.
@@ 3.2 Dual-layer EventSR SNN
Most previous works tend to increase the complexity of neural architectures for event stream super-resolution to achieve better performance (EF5; EF6). However, the baseline work has shown that for SNN-based event stream super-resolution, deeper networks may actually degrade performance (ES1), which supports our motivation to further simplify the network design.
@@
We propose a compact yet effective SNN-based event stream super-resolution network, simply named Dual-layer SNN in this paper. As shown in Figure 3 and Table 1, this network consists of two primary spiking convolution layers. The input and output of the network contain two channels, representing positive and negative event streams. The first layer extracts local spatiotemporal spike patterns and generates intermediate spike responses. The second layer upsamples the output of the first layer back to the original resolution. As shown in Figure 5, it also incorporates an upsampling bilinear-interpolated PSP bypass to preserve fine details from the low-resolution input for final reconstruction.
@@ 3.3 Dual-Forward Polarity-split Event Encoding
Recent studies have proposed using dual-branch networks to separately process positive and negative events (EF5; EF6), demonstrating that handling polarities separately leads to improved performance on event super-resolution. However, the additional network branches significantly increase the model’s weight.
@@ 3.5 Learnable Spatio-temporal Polarity-aware Loss
Inspired by previous work that performs the spatio-temporal learning (ES1) and weighted uncertainty loss (LearnableLoss), we extend the spatio-temporal only loss with polarity awareness and learnable adaptive weighting, namely Learnable Spatio-temporal Polarity-aware Loss (LearnSTPLoss), simply marked as ”w/L” in this paper. Given the predicted spike output $e_{\text{out}}$ and the ground truth spike stream $e_{\text{gt}}$, we have:
@@
Preserving time, space, and polarity information is essential to the event stream super-resolution task. When using a combined loss function, previous methods often rely on manually set hyperparameters to weight each component (ES1). However, such manual tuning lacks adaptability across different datasets or model architectures. To address this issue, we introduce a learnable weighting mechanism based on the logarithm of variance, enabling the weights of each loss term to be jointly balanced and optimized with the network parameters during training. This learnable loss formulation is expected to improve super-resolution accuracy without compromising the model’s lightweight design.
@@ 4.1 Experimental Settings
We evaluate our method on five datasets, as summarized in Table 3. The HR event streams correspond to the original event data in these datasets. To generate the LR event streams, we perform spatial downsampling by merging events within each 2×2 region using a stride of 2, which is shown in Figure 3.
Implementation Details.
We implement our SNN architecture using SlayerSNN (SlayerSNN), with a discretized simulation step size of 1 millisecond to update membrane potentials and determine spike firing. We train our model for 30 epochs using the Adam optimizer with an initial learning rate of 0.1. Table 3 provides the batch size for training each of the datasets. All experiments are conducted on an NVIDIA RTX 5090 GPU.
4.2 Evaluation Metrics
To quantitatively evaluate the performance of event stream super-resolution, referring to our baseline (ES1), we implement a Root Mean Squared Error (RMSE) metric that considers both temporal (MSET) and spatial (MSES) precision. The predicted and ground-truth event streams are represented as four-column arrays $(t,x,y,p)$ and discretized into 4D voxel tensors of shape (C=2,H,W,T)(C{=}2,H,W,T), where the two channels correspond to positive and negative polarities. In the formulations below, $(i,j)$ index pixel locations, $t$ denotes the temporal coordinate within the interval $T_{0},T_{1}]$, $N_{p}$ is the number of active pixels, $k$ indexes the temporal bins, and $N_{b}$ is the total number of bins.
Dataset Train/Test Sensor bs
N-MNIST (NMnist) 60k / 10k ATIS 64
CIFAR10-DVS (Cifar) 8.5k / 1.5k DVS128 8
ASL-DVS (ASL) 80.6k / 20.2k DAVIS240C 32
EventCameraD. (EventCameraDataset) 10.1k / 1.6k DAVIS240C 64
EventNFS-real (EF3) 74.2k / 10.5k DAVIS346 64
Table 3: Datasets and Batch Size (bs) information.
@@ 4.2 Evaluation Metrics
D-L .2621 .1292 .2280 97.34 .1781 .1002 .1473 69.92 .2250 .1048 .1975 99.39 .3165 .2593 .1697 89.06 .3180 .2160 .2329 87.46
D-L w/L .2607 .1277 .2272 97.87 .1772 .1017 .1451 70.17 .2250 .1038 .1982 99.42 .3130 .2626 .1675 89.82 .3105 .2148 .2237 87.65
U .2614 .1289 .2274 95.63 .1754 .0992 .1446 68.28 .2247 .1069 .1961 98.87 .3128 .2626 .1674 89.85 .3113 .2123 .2270 89.35
U w/L .2611 .1283 .2274 95.74 .1747 .0994 .1437 68.78 .2236 .1067 .1948 98.93 .3117 .2592 .1699 89.03 .3061 .2118 .2202 89.96
Table 4: Comparison of methods and baseline. Dual-Layer SNN and Ultralight SNN are marked as ”D-L” and ”U”. LearnSTPLoss is marked as ”w/L”. (The best; The second best.)
Temporal Consistency Error ($\mathrm{MSE}_{T}$).
The temporal error is computed by summing the squared differences of the voxel tensors across all spatiotemporal locations (ES1). Following our detailed definition, it is formally expressed as:
$\mathrm{MSE}_{T}=\frac{1}{N_{p}}\sum_{i,j}\int_{T_{0}}^{T_{1}}\left(\text{Spike}_{i,j}^{h}(t)-\text{Spike}_{i,j}^{gt}(t)\right)^{2}dt.$ (15)
Spatial Consistency Error ($\mathrm{MSE}_{S}$).
To evaluate local event distribution over time, the event stream is divided into non-overlapping time blocks. For each block, a Peri-Stimulus Time Histogram (PSTH) is computed by summing voxel counts along the time dimension. $\mathrm{MSE}_{S}$ is defined as:
$\mathrm{MSE}_{S}=\frac{1}{N_{p}}\sum_{k=1}^{N_{b}}\sum_{i,j}\left\|\text{PSTH}_{i,j}^{h}(k)-\text{PSTH}_{i,j}^{gt}(k)\right\|_{2}^{2}.$ (16)
Root Mean Squared Error ($\mathrm{RMSE}_{\mathrm{ST}}$).
@@
$\mathrm{RMSE}_{\mathrm{ST}}=\sqrt{\frac{1}{(T_{1}-T_{0})\cdot N_{p}}\left(\mathrm{MSE}_{T}+\mathrm{MSE}_{S}\right)}.$ (17)
Polarity Accuracy (PA).
@@ 4.3 Comparative Results
Method Params FLOPs Time
@@
Dual-Layer SNN w/L (ours) 464 (0.464K) 0.53G 1.49ms
Ultralight SNN w/L (ours) 232 (0.232K) 0.58G 1.55ms
Baseline (ES1) 1040 (1.04K) 1.10G 1.62ms
@@
SRFBN-esr (SF5-21) 2.1M (2100K) 39.5G 37.3ms
RLSP-esr (SF5-8) 1.2M (1200K) 23.1G
RSTT-esr (SF5-13) 3.8M (3800K) 22.3G 61.4ms
@@
EventZoom (EF3) 11.5M (11500K) 65.3G 17.4ms
RecEvSR (EF4) 1.8M (1800K) 2.80G 13.2ms
BMCNET (EF5) 2.6M (2600K) 35.35G
RMFNET (EF6) 3.0M (3000K) 8.73G 7.0ms
Table 5: Comparison of model size, computational cost, and inference time on EventNFS dataset. †: Ultralight SNN uses concurrent dual forwards. Parts of the data are from (EF5).
Figure 10: Visualization Results on Real-world Deployment.
@@
Referring Table 6, we evaluate our method on the image reconstruction task using the Event Camera Dataset and E2VID (E2VID). The results show that our method improves SSIM by 7% and reduces MSE by 4.48% compared to the baseline. Figure 9 visualizes the reconstruction, where zoomed-up regions reveal that our method preserves clearer edge details.
Additionally, we conduct classification tasks on N-MNIST, ASL-DVS, and CIFAR10-DVS datasets. Our method consistently improves the performance (See Appendix (appendix)). These results also demonstrate that higher-resolution event data leads to better performance on these tasks, further validating the significance of super-resolving event data.
Metric LR HR-GT (a) Baseline (b) (c)
@@
Table 6: Results of Image Reconstruction on Event Camera Dataset. (a) (EF1) (b) Dual-Layer SNN w/L (c) Ultralight SNN w/L
@@
Since event-to-frame methods evaluate only the spatial quality of their frame-based HR outputs, our method is not directly comparable at the event stream level. In the appendix, we provide further analytical comparisons in terms of spatiotemporal quality. We also deploy a DVXplorer-Lite event camera in a mobile scenario to collect real-world data and evaluate our method, demonstrating its robustness in performing 2× super-resolution on raw event streams. The visualization results can be found in Figure 10. (See Appendix (appendix) for more details)